OpenStudy (anonymous):

how far have you traveled if you drive 5 1/2 hours at 59mph?

OpenStudy (anonymous):

324.5 miles. You multiply 5.5 times 59.

OpenStudy (anonymous):

distance = rate x time = 59 mph * (5 1/2 hr) = 59 mph * (11 / 2 hr) = 649 / 2 miles = 324.5 miles or . . 325

OpenStudy (anonymous):

thankyou that really helped!